# sample.bib

### v4

@@ Abstract
Recently, the use of transformers in offline reinforcement learning has become a rapidly developing area. This is due to their ability to treat the agent’s trajectory in the environment as a sequence, thereby reducing the policy learning problem to sequence modeling. In environments where the agent’s decisions depend on past events (POMDPs), capturing both the event itself and the decision point in the context of the model is essential. However, the quadratic complexity of the attention mechanism limits the potential for context expansion. One solution to this problem is to enhance transformers with memory mechanisms. This paper proposes a Recurrent Action Transformer with Memory (RATE), a novel model architecture incorporating a recurrent memory mechanism designed to regulate information retention. To evaluate our model, we conducted extensive experiments on memory-intensive environments (ViZDoom-Two-Colors, T-Maze, Memory Maze, Minigrid.Memory), classic Atari games and MuJoCo control environments. The results show that using memory can significantly improve performance in memory-intensive environments while maintaining or improving results in classic environments. We hope our findings will stimulate research on memory mechanisms for transformers applicable to offline reinforcement learning.
@@ 1 Introduction
Figure 1: Recurrent Action Transformer with Memory.
@@
Transformers [vaswani2017attention], originally developed for natural language processing (NLP), perform well in Reinforcement Learning (RL) [li2023a, agarwal2023transformers]: online RL [parisotto2020stabilizing, lampinen2021towards, esslinger2022dtqn, zheng2022online, melo2022transformers, team2023human, pramanik2023recurrent], offline RL [chen2021decision, janner2021offline, lee2022multi, jiang2023efficient], and model-based RL [chen2022transdreamer, micheli2023transformers, robine2023transformerbased], including for solving credit assignment problems and working in memory-intensive environments [chen2021decision, ni2023when, grigsby2024amago], provided that the entire trajectory fits within the model context. Transformers struggle with long sequences due to quadratic attention complexity, limiting their use in episodic and continuing tasks. Various approaches attempt to increase context size  [dai2019transformerxl, zaheer2020big, beltagy2020longformer, choromanski2020rethinking, bulatov2022rmt, ding2023longnet], but such models may become unstable when training on long sequences [zhang2022opt] or to use a specific sparse attention mechanism which is unsuitable for non-NLP tasks [zaheer2020big, beltagy2020longformer, ding2023longnet]. Memory mechanisms offer a promising solution to consider past information without increasing context size. Our work explores memory in transformers for RL, inspired by NLP findings [bulatov2022rmt]. The RL setting differs from NLP in input sequence processing, requiring specialized encoders for observations, rewards, and actions, and are also characterized by significantly longer sequence lengths during inference.
@@
In RL, memory has 2 senses. One is using past information within an episode to make decisions [lampinen2021towards, ni2023when]. The other is transferring experience from one environment to another [melo2022transformers, team2023human, kang2023think], improving generalizability and solving Meta-RL [wang2016learning, duan2016rl] tasks. Our work focuses on the first sense [ni2023when]: using past information to make decisions within the same episode.
@@
In our work, we propose the Recurrent Action Transformer with Memory (RATEFigure 1), a model that uses memory mechanism: memory embeddings and Memory Retention Valve (MRV). We empirically show that memory embeddings effectively preserve information from previous steps, allowing the model to use past information when making decisions in the present. MRV is designed to control the process of updating memory embeddings and prevent the loss of important information when processing long sequences. We conduct thorough experiments on the memory-intensive environments [memup, pasukonis2022memmaze, MinigridMiniworld23, ni2023when] and on the standard RL benchmarks [bellemare2013arcade, d4rl]. We also study the impact of memory on the performance of the proposed model.
@@
    We propose the Recurrent Action Transformer with Memory (RATE), a transformer model that uses a memory mechanism.
@@
    We showed that RATE significantly outperforms counterparts without memory mechanism in memory-intensive environments: ViZDoom-Two-Colors, MemoryMaze, Minigrid.Memory and T-Maze.
@@
    We showed that RATE demonstrates better or comparable results in classic Atari games and MuJoCo control tasks. These tasks do not require the use of memory explicitly, demonstrating that the proposed mechanism is suitable for different types of tasks and emphasizing its universality.
@@ 2.1 Offline Reinforcement Learning
In reinforcement learning [sutton2018reinforcement], we assume that the task can be described by a Markov Decision Process as a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}\rangle$. The process consists of states $s\in\mathcal{S}$, actions $a\in\mathcal{A}$, a state transition function $\mathcal{P}(s^{\prime}|s,a)$, and an immediate reward function $r=\mathcal{R}(s,a)$. The states are assumed to have the Markov Property, that is $\mathbb{P}(s_{t+1}|s_{t})=\mathbb{P}(s_{t+1}|s_{1},\dots,s_{t})$. Given a timestep $t$, we use $r_{t}=R(s_{t},a_{t})$ to denote the immediate reward that the agent receives at state $s_{t}$ performing action $a_{t}$ at that timestep. We describe trajectory $\tau$ of length $T$ as a sequence of states $s_{i}$, actions $a_{i}$, and immediate rewards $r_{i}$: $\tau=(s_{0},a_{0},r_{0},s_{1},a_{1},r_{1},\dots,s_{T},a_{T},r_{T})$. We denote return-to-go [chen2021decision] $R_{t}$ of trajectory $\tau$ at the timestep $t$ as a sum of future rewards from the timestep $t$ to the end of trajectory: $R_{t}=\sum^{T}_{t^{\prime}=t}r_{t^{\prime}}$. The goal of a reinforcement learning agent is to learn policy $\pi$ that maximizes the expected return. In online RL, the trajectories used to train an agent are obtained iteratively as the agent interacts with the environment. In offline RL, the agent does not interact with the environment during training. A fixed set of trajectories collected by an arbitrary policy is used for training. Although such a setting is more difficult because it does not allow additional exploration of the environment or generation of new trajectories, it is preferable for tasks where interaction with the environment is costly or risky, such as in robotics.
@@ 2.2 Partially Observable Markov Decision Process
In the real world, there are frequent situations where the full state of the environment is not available to the agent, so the Markov Property is violated, and the agent is said to receive observations instead of state as input. In this case, observations are no longer sufficient statistics of the past to make a decision in the current step. An example would be a robot navigating an environment based on a camera image or a situation in which a decision must be made based on information from the past that is not available in the current observation. A Partially Observable Markov Decision Process (POMDP) is used in such cases. POMDP is a generalization of the Markov Decision Process (MDP) and is written as $\langle\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{P},\mathcal{R},\mathcal{Z}\rangle$, where $o\in\mathcal{O}$ – observations, $s\in\mathcal{S}$ – states, $a\in\mathcal{A}$ – actions, $r=\mathcal{R}(s,a)$ – immediate reward function, $\mathcal{P}(s^{\prime}|s,a)$ – state transition function and $\mathcal{Z}$ is an observation function $\mathcal{Z}^{a}_{s^{\prime}o}=P(O=o_{t+1}|S_{t+1}=s^{\prime},A_{t}=a)$. To work successfully in such environments, mechanisms such as memory are needed to allow the use of information from the past. In our work, we propose an approach of adding memory to the agent in the form of memory embeddings and MRV. We consider the formulation of an offline model-free RL, where learning is formulated as a sequence modeling problem [chen2021decision].
@@ 3 Recurrent Action Transformer with Memory
In this paper, we introduce a new architecture – Recurrent Action Transformer with Memory (RATE) in which we utilized both recurrently trained memory embeddings [bulatov2022rmt] and the preservation of previous hidden states [dai2019transformerxl]. A Memory Retention Valve (MRV) inside our model allows important information to be carried through sequences many times longer the length of the context. The architecture of RATE is shown in Figure 1.
We used the methodology outlined in the Decision Transformer paper [chen2021decision] to represent the modeled trajectory. The trajectory is represented as a sequence of triplets $(R,o,a)$, where $R$ denotes the return-to-go, $o$ represents the observation, and $a$ signifies the action. The return-to-go $R_{t}$ at time step $t$ corresponds to $R_{t}=\sum^{T}_{t^{\prime}=t}r_{t^{\prime}}$, and the $R_{t+1}$ at time step $t+1$ is obtained from $R_{t}$ by subtracting the immediate reward $r_{t}$ at time step $t$: $R_{t+1}=\sum^{T}_{t^{\prime}=t+1}r_{t^{\prime}}=R_{t}-r_{t}$. Each component is encoded using the appropriate encoder (Table 6). The memory embeddings, which are model parameters, are inserted at the beginning and end of each segment in the input sequence. This embedded sequence serves as the input for the transformer. The resulting embeddings, obtained in the transformer’s output, correspond to observation vectors that are further employed with linear projection to predict the agent’s action in the respective frame. Memory embedding is a subsequence of $M$ vectors, where $M$ is a hyperparameter. These memory embedding subsequences are the same at the beginning and the end of the sequence in the current segment.
During training, we split the trajectories into $N$ segments of length $K$. Thus, instead of the notion of a context $K$, we use the notion of an effective context $K_{eff}=N\times K$ [bulatov2022rmt], which consists of $N\times K$ environment steps. We initialize the memory embeddings $M_{0}$ with normally distributed random trainable memory vectors. Following a forward pass, we retrieve a new memory embeddings $\tilde{M}_{n+1}$ from the MRV part of RATE model, which is designed to control the process of forgetting important information. To avoid computational instability, we detach the new memory embeddings. However, the gradient from the next segment can flow to the previous segment through this memory embeddings. We used lower triangular attention masks to attend to tokens within the trajectory sequence. In the case of memory segments, a full attention mask was utilized. Subsequent segments entail updating the memory vectors based on the extracted representations of the suffix memory vectors at the top layer of the transformer. The MRV block also uses a full attention mask when processing memory embeddings.
A sequence of alternately concatenated sequences of memory embeddings $M_{n}$, segment, and memory embeddings $M_{n}$ is fed to the RATE input on the segment $n\in[0,N-1]$. As output, we get a sequence of predicted in auto-regressive manner agent’s actions $\hat{a}_{K}$ of length $K$, used to compute loss, and the updated values of memory embeddings $\hat{M}_{n+1}$, that are input to the RATE input together with the next segment:
$\hat{a}_{K},\tilde{M}_{n+1}=RATE(\text{concat}(M_{n},\text{segment},M_{n}))$ (1)
In more detail, the architecture is as follows:
First, we feed a concatenated sequence consisting of memory embeddings and a segment to the input of the Causal Transformer:
$\hat{a}_{K},M_{n+1}=Transformer(\text{concat}(M_{n},\text{segment},M_{n}))$ (2)
We obtain the predicted actions $\hat{a}_{K}$ used to compute loss, and the updated memory embeddings $M_{n+1}$. We then feed these new memory embeddings $M_{n+1}$ together with those $M_{n}$ fed to the Causal Transformer input to the MRV:
$\tilde{M}_{n+1}=MRV(M_{n},M_{n+1})$ (3)
Thus, we get the final memory embeddings $\tilde{M}_{n+1}$, which will be used when processing the next segment. In more detail, inside MRV, we first compute Query, Key and Value matrices using cross-attention:
$Q_{h}=M_{n}W_{q,h}^{T},K_{h}=M_{n+1}W_{k,h}^{T},V_{h}=M_{n+1}W_{v,h}^{T}$ (4)
Then we use $Softmax()$ to process raw attention scores:
$\tilde{M}_{n+1}^{h}=Softmax(\frac{Q_{h}K_{h}^{T}}{\sqrt{d_{k}}})V_{h}$ (5)
Finally, we combine the matrices obtained for each attention head:
$\tilde{M}_{n+1}=Linear(\text{concat}(\tilde{M}_{n+1}^{0},...,\tilde{M}_{n+1}^{h}))$ (6)
Here $h$ is the head number of the MRV attention heads. In MRV, we use full attention masks.
During inference, actions are predicted in an autoregressive manner: $t_{0}:(R_{0},o_{0})\rightarrow RATE\rightarrow a_{0}$; $t_{1}:(R_{0},o_{0},a_{0},R_{1},o_{1})\rightarrow RATE\rightarrow a_{1}\ldots$

### v5

# Recurrent Action Transformer with Memory

@@ Abstract
Recently, the use of transformers in offline reinforcement learning has become a rapidly developing area. This is due to their ability to treat the agent’s trajectory in the environment as a sequence, thereby reducing the policy learning problem to sequence modeling. In environments where the agent’s decisions depend on past events (POMDPs), it is essential to capture both the event itself and the decision point in the context of the model. However, the quadratic complexity of the attention mechanism limits the potential for context expansion. One solution to this problem is to extend transformers with memory mechanisms. This paper proposes a Recurrent Action Transformer with Memory (RATE), a novel model architecture that incorporates a recurrent memory mechanism designed to regulate information retention. To evaluate our model, we conducted extensive experiments on memory-intensive environments (ViZDoom-Two-Colors, T-Maze, Memory Maze, Minigrid-Memory), classic Atari games, and MuJoCo control environments. The results show that using memory can significantly improve performance in memory-intensive environments, while maintaining or improving results in classic environments. We believe that our results will stimulate research on memory mechanisms for transformers applicable to offline reinforcement learning.
@@ 1 Introduction
Figure 1: Recurrent Action Transformer with Memory (RATE). $R$ – returns-to-go, $o$ – observations, $a$ – actions, $M_{n}$ – segment’s $S_{n}$ memory embeddings.
@@
Transformers (Vaswani et al., 2017), originally developed for Natural Language Processing (NLP), perform well in Reinforcement Learning (RL) (Agarwal et al., 2023; Li et al., 2023): online RL (Parisotto et al., 2020; Esslinger et al., 2022; Melo, 2022; Team et al., 2023), offline RL (Chen et al., 2021; Lee et al., 2022; Jiang et al., 2023), and model-based RL (Chen et al., 2022; Micheli et al., 2023; Robine et al., 2023), including solving the credit assignment problem and working in memory-intensive environments (Chen et al., 2021; Ni et al., 2023; Grigsby et al., 2024), provided that the entire trajectory fits within the model context. However, transformers struggle with long sequences due to quadratic attention complexity, limiting their use in long inference tasks. Several approaches attempt to increase the context size (Dai et al., 2019; Bulatov et al., 2022; Ding et al., 2023), but such models may become unstable when trained on long sequences (Zhang et al., 2022), or use a specific sparse attention mechanism that is unsuitable for non-NLP tasks (Beltagy et al., 2020; Zaheer et al., 2020; Ding et al., 2023). Memory mechanisms offer a promising solution to account for past information without increasing context size. Our work explores memory in transformers for RL, inspired by NLP results (Dai et al., 2019; Bulatov et al., 2022). The RL setting differs from NLP in the processing of input sequences, requiring specialized encoders for observations, rewards, and actions, and is also characterized by significant sparsity in some tasks.
@@
In RL memory has two senses. One is using past information within an episode to make decisions (Lampinen et al., 2021; Ni et al., 2023). The other is transferring experience from one environment to another (Melo, 2022; Kang et al., 2023; Team et al., 2023), improving generalizability, sample efficiency, and solving Meta-RL (Duan et al., 2016; Wang et al., 2016) tasks. Our work focuses on the first case (Ni et al., 2023): using past information to make decisions within the same episode.
@@
In this paper, we propose the Recurrent Action Transformer with Memory (RATEFigure 1), a model that uses several memory mechanisms: memory embeddings, caching of previous hidden states, and Memory Retention Valve (MRV). We empirically show that memory mechanisms effectively preserve information from previous steps, allowing the model to use past information when making decisions in the present. MRV is designed to control the process of updating memory embeddings and prevent the loss of important information when processing long sequences, thus enabling the processing of highly sparse tasks. To evaluate the memory mechanisms, we perform extensive experiments in memory-intensive environments: ViZDoom-Two-Colors (Sorokin et al., 2022), Memory Maze (Pasukonis et al., 2022), Minigrid-Memory (Chevalier-Boisvert et al., 2023), and Passive T-Maze (Ni et al., 2023), as well as on standard RL benchmarks: Atari (Bellemare et al., 2013) and MuJoCo (Fu et al., 2021). We also study the impact of memory on the performance of the proposed model.
The proposed model interpolates and extrapolates well outside the transformer context, is able to retain important information for a long time when operating in highly sparse environments, and allows to compensate the effect of bias in the training data.
@@
    We propose the Recurrent Action Transformer with Memory (RATE), a transformer model for offline RL that makes use of memory mechanisms: memory embeddings, caching of hidden states, and the Memory Retention Valve (MRV). The proposed MRV is based on the cross-attention architecture and is designed to prevent information loss from memory embeddings and significantly improve the performance of RATE in memory-intensive environments with sparse structure (see section 3).
@@
    We show that RATE significantly outperforms strong baselines with and without memory mechanisms in memory-intensive environments, including ViZDoom Two-Colors, Memory Maze, Minigrid-Memory, and T-Maze (see subsection 4.2).
@@
    We demonstrate that RATE achieves better or comparable results in classic Atari games and MuJoCo control tasks, demonstrating that the proposed model is suitable for different types of tasks and emphasizing its universality (see subsection 4.2).
@@ 2.1 Offline Reinforcement Learning
In RL (Sutton & Barto, 2018), we assume that the task can be described by a Markov Decision Process (MDP) as a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}\rangle$. The process consists of states $s\in\mathcal{S}$, actions $a\in\mathcal{A}$, a state transition function $\mathcal{P}(s^{\prime}|s,a)$, and an immediate reward function $r=\mathcal{R}(s,a)$. The states are assumed to have the Markov Property, that is $\mathbb{P}(s_{t+1}|s_{t})=\mathbb{P}(s_{t+1}|s_{1},\dots,s_{t})$. Given a timestep $t$, we use $r_{t}=R(s_{t},a_{t})$ to denote the immediate reward that the agent receives at state $s_{t}$ performing action $a_{t}$ at that timestep. We describe trajectory $\tau$ of length $T$ as a sequence of states $s_{i}$, actions $a_{i}$, and immediate rewards $r_{i}$: $\tau=(s_{0},a_{0},r_{0},s_{1},a_{1},r_{1},\dots,s_{T-1},a_{T-1},r_{T-1})$. We denote return-to-go (Chen et al., 2021) $R_{t}$ of trajectory $\tau$ at the timestep $t$ as a sum of future rewards from the timestep $t$ to the end of trajectory: $R_{t}=\sum^{T-1}_{t^{\prime}=t}r_{t^{\prime}}$. The goal of a RL agent is to learn policy $\pi$ that maximizes the expected return. In online RL, the trajectories used to train an agent are obtained iteratively as the agent interacts with the environment. In offline RL, the agent does not interact with the environment during training. A fixed set of trajectories collected by an arbitrary policy is used for training. Although such a setting is more difficult because it does not allow additional exploration of the environment or generation of new trajectories, it is preferable for tasks where interaction with the environment is costly or risky, such as in robotics.
@@ 2.2 Partially Observable Markov Decision Process
In the real world, there are frequent situations where the full state of the environment is not available to the agent, so the Markov Property is violated, and the agent is said to receive observations instead of state as input. In this case, observations are no longer sufficient statistics of the past to make a decision in the current step. An example would be a robot navigating an environment based on a camera image or a situation in which a decision must be made based on information from the past that is not available in the current observation. A Partially Observable Markov Decision Process (POMDP) is used in such cases. POMDP is a generalization of the MDP and is written as $\langle\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{P},\mathcal{R},\mathcal{Z}\rangle$, where $o\in\mathcal{O}$ – observations, $s\in\mathcal{S}$ – states, $a\in\mathcal{A}$ – actions, $r=\mathcal{R}(s,a)$ – immediate reward function, $\mathcal{P}(s^{\prime}|s,a)$ – state transition function and $\mathcal{Z}$ is an observation function $\mathcal{Z}^{a}_{s^{\prime}o}=P(O=o_{t+1}|S_{t+1}=s^{\prime},A_{t}=a)$. To work successfully in such environments, mechanisms such as memory are needed to allow the use of information from the past (Parisotto et al., 2020; Lampinen et al., 2021; Esslinger et al., 2022). In our work, we propose an approach of adding memory to the agent in the form of memory embeddings, caching of hidden states, and MRV. In this paper we consider the setting of an offline model-free RL, where learning is formulated as a sequence modeling problem (Chen et al., 2021).
@@ 3 Recurrent Action Transformer with Memory
Algorithm 1 Recurrent Action Transformer with Memory
Input: $R\in\mathbb{R}^{T},o\in\mathbb{R}^{d_{o}\times T},a\in\mathbb{R}^{T}$
Parameters: $M\in\mathbb{R}^{m\times d}$
1:  $\tilde{R}\in\mathbb{R}^{T\times d}\leftarrow\texttt{Encoder}_{R}(R)$ $\ \tilde{o}\in\mathbb{R}^{T\times d}\leftarrow\texttt{Encoder}_{o}(o)$ $\ \tilde{a}\in\mathbb{R}^{T\times d}\leftarrow\texttt{Encoder}_{a}(a)$
2:  $\tau_{0:T-1}\leftarrow\{(\tilde{R}_{0},\tilde{o}_{0},\tilde{a}_{0}),\ldots,(\tilde{R}_{T-1},\tilde{o}_{T-1},\tilde{a}_{T-1})\}$
3:  $M\leftarrow M_{0}\sim\mathcal{N}(0,1)$
4:  for n $\textbf{in range }[0,T//K-1]$ do
5:     $S_{n}\in\mathbb{R}^{3K\times d}\leftarrow\tau_{n\times K:(n+1)\times K}$
6:     $\tilde{S}_{n}\in\mathbb{R}^{(3K+2m)\times d}\leftarrow\texttt{concat}(M_{n},S_{n},M_{n})$
7:     $\hat{a}_{n},M_{n+1}\leftarrow\texttt{Transformer}(\tilde{S}_{n})$
8:     $M_{n+1}\leftarrow\texttt{MRV}(M_{n},M_{n+1})$Output: $\hat{a}_{n}\rightarrow\mathcal{L}(a_{n},\hat{a}_{n}),\ M_{n+1}$
9:  end for
Algorithm 2 Memory Retention Valve
Input: $M_{n},M_{n+1}\in\mathbb{R}^{m\times d}$
Parameters: $\mathbf{W}_{Q}^{h},\mathbf{W}_{K}^{h},\mathbf{W}_{V}^{h}\in\mathbb{R}^{d_{h}\times d},\mathbf{W}_{M}\in\mathbb{R}^{d\times d}$
1:  $\mathbf{Q}_{h}\leftarrow M_{n}\mathbf{W}_{Q}^{h\ T}$
2:  $\mathbf{K}_{h}\leftarrow M_{n+1}\mathbf{W}_{K}^{h\ T}$
3:  $\mathbf{V}_{h}\leftarrow M_{n+1}\mathbf{W}_{V}^{h\ T}$
4:  $M_{n+1}^{h}\leftarrow\texttt{softmax}(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}^{T}}{\sqrt{d}})\mathbf{V}_{h}$
5:  $M_{n+1}\leftarrow\texttt{concat}(M_{n+1}^{0},\ldots,M_{n+1}^{h})$
6:  $M_{n+1}\leftarrow M_{n+1}\mathbf{W}_{M}^{T}$
Output: $M_{n+1}$
@@
In this paper, we introduce a new architecture, Recurrent Action Transformer with Memory (RATE), in which we utilized recurrently trained memory embeddings (Bulatov et al., 2022) and caching of previous hidden states (Dai et al., 2019) to add memory, and Memory Retention Valve (MRV) to control information leakage from memory embeddings, allowing sparse sequences to be processed. The architecture of RATE is shown in Figure 1.
The RATE scheme is outlined in Algorithm 1. In the initial phase, we obtain embeddings $\tilde{R}$, $\tilde{o}$ and $\tilde{a}$ from the returns-to-go $R$, observations $o$, and actions $a$, respectively, using the corresponding encoders from Table 10. We then generate a trajectory $\tau_{0:T-1}$ consisting of triplets of these embeddingsx according to the technique described in the Decision Transformer (DT) (Chen et al., 2021) paper (Algorithm 3).
Next, the trajectory $\tau_{0:T-1}$ is divided into $N=T//K$ segments $S_{n}\in\mathbb{R}^{3K\times d},\ n\in[0,N-1]$, each consisting of $K$ triplets, where $K$ is the context length and $d$ is the model dimension. To each segment $S_{n}$, memory embeddings $M_{n}\in\mathbb{R}^{m\times d}$ are concatenated at the beginning and at the end, forming a $\tilde{S}_{n}\in\mathbb{R}^{(3K+2m)\times d}$, where $m$ is the number of memory embeddings. These segments $\tilde{S}_{n}$ are then fed into the transformer and the output is the predicted actions $\hat{a}_{n}=\hat{a}_{n\times K:(n+1)\times K}$, which are used to compute loss, and new memory embeddings $M_{n+1}$, which are then processed by the MRV block and transmitted to the next segment $S_{n+1}$.
The MRV scheme is presented in Algorithm 2 and is based on the cross-attention basis. After the transformer processes the segment $S_{n}$, the previous memory embeddings $M_{n}$ and the new memory embeddings $M_{n+1}$ obtained at the output of the transformer are fed to the input of the MRV block. Next, $M_{n}$ are multiplied by the query matrix $\mathbf{W}_{Q}^{h}$, and $M_{n+1}$ are multiplied by the key $\mathbf{W}_{K}^{h}$ and value $\mathbf{W}_{V}^{h}$ matrix for each of the attention heads $h$. Then the attention scores are calculated using the $\texttt{softmax}()$ function, the results for each of all attention heads are concatenated and after linear transformation using the matrix $\mathbf{W}_{M}^{T}$ the final memory tokens $M_{n+1}$ are obtained at the output, which is passed to the next segment.
Unlike the DT learning process, where random fragments of length $K$ are cut from trajectories, we process segments of length $K$ sequentially, which allows us to capture all information in the processed trajectory. Thus, using memory mechanisms, we are able to process sequences of length $K_{eff}=K\times N$, where $K_{eff}$ is the effective context (Bulatov et al., 2022).